# On the hierarchical Bayesian modelling of frequency response functions

### v1

@@ Abstract
Population-based structural health monitoring (PBSHM) aims to share valuable information among members of a population, such as normal- and damage-condition data, to improve inferences regarding the health states of the members. Even when the population is comprised of nominally-identical structures, benign variations among the members will exist as a result of slight differences in material properties, geometry, boundary conditions, or environmental effects (e.g., temperature changes). These discrepancies can affect modal properties and present as changes in the characteristics of the resonance peaks of the frequency response function (FRF). Many SHM strategies depend on monitoring the dynamic properties of structures, so benign variations can be challenging for the practical implementation of these systems. Another common challenge with vibration-based SHM is data loss, which may result from transmission issues, sensor failure, a sample-rate mismatch between sensors, and other causes. Missing data in the time domain will result in decreased resolution in the frequency domain, which can impair dynamic characterisation. The hierarchical Bayesian approach provides a useful modelling structure for PBSHM, because statistical distributions at the population and individual (or domain) level are learnt simultaneously to bolster statistical strength among the parameters. As a result, variance is reduced among the parameter estimates, particularly when data are limited. In this paper, combined probabilistic FRF models are developed for a small population of nominally-identical helicopter blades under varying temperature conditions, using a hierarchical Bayesian structure. These models address critical challenges in SHM, by accommodating benign variations that present as differences in the underlying dynamics, while also considering (and utilising), the similarities among the blades.
@@
Population-based structural health monitoring (PBSHM); Hierarchical Bayes; Multilevel models; Uncertainty quantification; Repeatability
[…]
@@ 1 Introduction
The current work is focussed on developing population-based structural health monitoring (PBSHM) techniques, for a group of nominally-identical structures (i.e., a homogeneous population L.A. Bull et al. (2021, 2019)), with consideration for the effects of temperature changes, to improve understanding of the variability in the health states of the population and the individual members. (In contrast, a heterogeneous population Gosliga et al. (2021); P.A. Gardner et al. (2021) will have greater discrepancies among the members, such as different suspension bridge designs, and may require further processing such as domain adaptation Gardner et al. (2020); P. Gardner et al. (2020, 2022); J. Poole et al. (2022)). Even among nominally-identical structures, variations caused by manufacturing differences, ageing parts, and changes in testing conditions can introduce uncertainty in the underlying dynamics. Cawley et al. P. Cawley et al. (1985) performed vibration testing on filament-wound carbon fibre-reinforced plastic (CFRP) tubes. Of the 18 tested, six tubes were considered ‘normal’ (i.e., having the same microstructure, with a ±$\pm 45^{\circ}$ fibre winding angle) P. Cawley et al. (1985). They found that the first and second natural frequencies of the ‘normal’ tubes varied by as much as 4% P. Cawley et al. (1985). When the remaining 12 tubes were also considered (which had intentional defects in their microstructure, such as slightly misaligned fibres, changes in volume fraction, etc.), the natural frequencies varied by as much as 18% P. Cawley et al. (1985). Changes in bolt tightness can similarly cause variations in natural frequency. Zhu et al. W.D. Zhu and He (2013) performed modal testing on an aluminium three-bay space-frame structure with bolted joints, before and after manually loosening several bolts to hand tight. They found that natural frequencies changed by as much as 8%, depending on the mode considered W.D. Zhu and He (2013). Global variations, such as changes in ambient temperature, can also affect modal properties. For example, increased temperature may reduce stiffness (depending on the specific material properties of the structure, and the duration/temperature of exposure), which may decrease natural frequency. Colakoglu M. Colakoglu (2008) tested a polyethylene fibre composite beam from ${}^{\circ}$ to ${}^{\circ}$, and found that the first natural frequency decreased by 12% over the measured temperature range. In addition, for the same polyethylene beam, Colakoglu M. Colakoglu (2008) found that modal damping increased with rising temperature. They concluded that the relationships between temperature and natural frequency, and between temperature and damping, appeared to be functional and monotonic for the polyethylene beam, over the temperature range considered M. Colakoglu (2008). Similar effects of temperature on stiffness/natural frequency and damping have been noted for a magnesium alloy H. Watanabe et al. (2004) and a composite honeycomb structure Y. Bai et al. (2018). Accounting for these benign fluctuations is important for the practical implementation and generalisation of SHM technologies, as features commonly used for damage identification may be sensitive to harmless changes as well as damage Worden et al. (2002); Sohn (2007); Alampalli (1998); Cawley (1997). (Indeed, it is well known that structural damage can reduce stiffness, often manifesting as a reduction in natural frequency. Benign variations can then either mimic or mask damage, depending on whether they exhibit a stiffening or softening effect Alampalli (1998); Cawley (1997); Sohn (2007). Similar effects have been noted with damping, and multiple studies have shown that damping tends to increase with damage, particularly with crack growth Cao et al. (2017).) Data scarcity and sparsity present additional challenges for SHM systems that rely on machine learning. Data scarcity refers to incomplete information regarding the damage- or normal-condition states of structures, particularly those newly in operation, and can impair model training and development. Likewise, sensing networks are prone to data loss (causing sparse data), because of sensor failure caused by harsh environmental conditions or insufficient maintenance. Transmission issues make wireless-sensing networks particularly susceptible to loss, and can be caused by large transmission distances between the sensors and base station J. Pei et al. (2005), software/hardware problems J. Meyer et al. (2010), and other issues such as weather changes, interference from nearby devices, or installation difficulties Y. Bao et al. (2013). Furthermore, modern systems that produce large amounts of high-resolution data can suffer losses resulting from a data transfer bottleneck A. Malekloo et al. (2022). Significant losses higher than 30% have been reported J. Meyer et al. (2010); Y. Bao et al. (2013); J. Lynch and Loh (2006), and a 0.38% data loss was found to have similar effects on power spectral density (PSD) as 5% additive noise T. Nagayama et al. (2007). Differences in sample rate among sensors could have a similar presentation, with the data captured at a lower rate seemingly missing data relative to that captured at a higher rate. Again, PBSHM addresses these scarcity/sparsity issues via knowledge transfer among similar structures, so that data-rich members can support those with more limited information. Homogenous populations can be represented using a general model, called a population form. The form attempts to capture the ‘essential nature’ of the population and benign variations among the members L.A. Bull et al. (2021). The form was first introduced in L.A. Bull et al. (2021, 2019), where a conventional or single Gaussian process (GP) was applied to frequency response functions (FRFs), to develop a representation for a nominally-identical population of eight degree-of-freedom (DOF) systems. Then, to accommodate greater differences among the nominally-identical members, an overlapping mixture of Gaussian processes (OMGP) Lázaro-Gredilla et al. (2012), was used in L.A. Bull et al. (2021, 2021), to infer multivalued wind-turbine power-curve data, with unsupervised categorisation of the data. The OMGP approach Lázaro-Gredilla et al. (2012) was again used in T.A. Dardeno et al. (2022) to develop a population form for real and imaginary FRFs, obtained from four nominally-identical, full-scale, composite helicopter blades. Recently, hierarchical modelling was used to improve the predictive capability of simulated L.A. Bull et al. (2022) and in-service L.A. Bull et al. (2023a) truck-fleet hazard models, and wind-turbine power curves L.A. Bull et al. (2023a). Specifically, the work presented in L.A. Bull et al. (2022, 2023a) used partial pooling, where data groups are treated as belonging to different domains, and each domain can be considered a realisation from global or population-level distributions. (This approach is in contrast to complete pooling, where all data are considered as belonging to a single domain, or no pooling, where each domain is treated as fully independent of the other domains). It was shown that when populations of structures were allowed to share correlated information, model uncertainty was reduced L.A. Bull et al. (2022, 2023a). In addition, domains with incomplete data were able to borrow statistical strength from data-rich groups L.A. Bull et al. (2022, 2023a). In L.A. Bull et al. (2023b), multilevel modelling with partial pooling was used to learn a 2D map of arrival times for waves propagating through a complex plate geometry, via GP regression, using a series of acoustic-emission experiments performed on the same plate but with differing experimental designs. Domain expertise regarding the positive gradient for the expected intertask functions was encoded into the model, via a linear mean function and appropriate priors L.A. Bull et al. (2023b); M.R. Jones et al. (2023). GP kernel hyperparameters were learnt at the domain level, to allow variation in the response surface among the different tests, and were correlated via a higher-level sampling distribution L.A. Bull et al. (2023b).
@@ 1.1 Research aims of the current work
The current work addresses critical SHM challenges, including uncertainties in the underlying dynamics of structures, and data-scarcity/sparsity issues, which can impair generalisation of SHM technologies. In line with the population forms developed in L.A. Bull et al. (2021, 2019, 2021); T.A. Dardeno et al. (2022); L.A. Bull et al. (2022, 2023a), this paper presents generalised, probabilistic FRF models that were developed for the helicopter blades in T.A. Dardeno et al. (2022), using a hierarchical Bayesian structure. Two case studies are presented. The first case used FRFs collected from all four blades, at ambient laboratory temperature, with variations among the blades resulting from manufacturing differences (e.g., small discrepancies in material properties and geometry) and boundary conditions. Limited training data that did not fully characterise the resonance peaks were taken from two of the FRFs, while sufficient training data were taken from the remaining two FRFs, so that information could be shared with the data-poor domains via shared distributions over the parameters. This situation is representative of incomplete data in the time domain, which would reduce the number of spectral lines in the frequency domain. Independent models were generated for comparison, to visualise the variance reduction from the combined model. The second case considered vibration data from one of the helicopter blades collected at various temperatures in an environmental chamber. A probabilistic FRF model was again developed using a hierarchical approach with partial pooling. Functional relationships between temperature and natural frequency, and between temperature and damping, were approximated via Taylor series expansion, and polynomial coefficients were learnt at the population level. A subset of temperature-varied FRFs were used to train the model. Population-level modal parameters and polynomial coefficients were then used to generalise to temperature-varied FRFs not used to train the model, and model accuracy was evaluated by comparing the results to FRFs computed via measured vibration data.
@@ 2 Related work
Most literature related to population-level monitoring of engineering structures is focussed on transfer learning, which aims to improve predictions in a target domain given a source domain with more complete information. Some cases have involved fine-tuning the classifiers and weights of a pre-trained convolutional neural network (CNN) according to a new dataset. For example, CNNs have been used to detect cracks, as in R-T. Wu et al. (2019); Dorafshan et al. (2018); Gao and K.M. Mosalam (2018); Jang et al. (2019), and other defects (e.g., corrosion, staining), as in R-T. Wu et al. (2019); Perez et al. (2019). Other cases have used domain adaptation (DA), a subcategory of transfer learning, where the source and target domains have the same feature space, and the target domain is mapped onto a shared space. Predictions are then made from a single model. For example, domain-adversarial neural networks have been implemented for condition monitoring of a fleet of power plants Michau and Fink (2019), and for fault detection in a gearbox and three-story structure Ozdagli and Koutsoukos (2020). A kernelised linear projection approach to DA has also been studied, between simulated source and target structures Gardner et al. (2020); P. Gardner et al. (2020), and between a simulated source structure and experimental target structure P. Gardner et al. (2022). DA has also been used to transfer damage detectors across systems using vibration data, in a group of tail planes L.A. Bull et al. (2021). Lastly, statistic alignment has been shown to improve the performance of established DA methods, by making them more robust to class imbalance and data-sparsity issues J. Poole et al. (2022). Population-level modelling can also be viewed in the context of multitask learning (MTL), where multiple tasks are solved simultaneously, while considering similarities (and differences) across tasks. In MTL, a combined inference allows sharing of correlated information among domain-specific models, to improve accuracy in data-poor domains L.A. Bull et al. (2023a); Sun et al. (2021). In the context of modelling engineering infrastructure, MTL has been used as part of a multioutput Gaussian process (GP) regression to learn correlations between data obtained at adjacent sensors from the Canton Tower, to reconstruct missing information H-P. Wan and Y-Q. Ni (2019). (Note that using MTL in this context differs from the current work. In H-P. Wan and Y-Q. Ni (2019), missing time-domain data from temperature and acceleration sensors were reconstructed using MTL, and all data were sourced from a single structure. The current work considers frequency-domain data from multiple structures, as in the first case, and incorporates the relationships between temperature and domain-specific tasks into the model structure, as in the second case.) A similar approach was used in Y. Li et al. (2021), where GP regression assisted with missing data recovery from a faulty sensor, by capturing the correlation among the remaining sensors on a hydroelectric dam. Likewise, in Seshadri et al. (2022), GPs were used to transfer information between spatial temperature or pressure profiles at axial stations within an aeroengine. Hierarchical Bayesian modelling is an MTL framework, where (lower-level) domain-specific tasks are correlated by conditioning on (higher or population-level), shared variables. The technique was first introduced in an SHM context in Huang and J.L. Beck (2015); Huang et al. (2019), where a multilevel hierarchical Bayesian model was developed to identify damage on a single structure, by inferring stiffness loss from changes in modal parameters, given noisy, incomplete modal data. (Note that this approach differs from the current work. In Huang and J.L. Beck (2015); Huang et al. (2019), modal parameters were learnt for different damage scenarios using a series of coupled linear regressions to approximate the eigenvalue problem. In Huang et al. (2019), the similarities between acceleration responses at adjacent sensors were exploited for data-loss recovery, by modelling the measured signal as a linear combination of basis functions. Both cases considered data from a single structure. The work presented in the current paper, in the first case study, models the entire (band-limited) FRF from multiple structures, using a likelihood function based on modal parameters. The second case presented in the current work models the entire FRF using measurements at various temperatures from a single structure, and incorporates functional relationships between temperature and the modal parameters, enabling predictions at ‘unseen’ temperature conditions.) Hierarchical Bayesian models were again used in D. Di Francesco et al. (2021), to estimate corrosion rates via data from multiple sensors on the same test structure, to support decision-making in the absence of complete data. In Papadimas and Dodwell (2021), hierarchical Bayesian models were developed using strain measurements from multiple tensile tests to inform the material properties of the samples, in a manner that considered the inherent variability in material properties among the samples and uncertainties related to experimental repeatability. Likewise, in Dhada et al. (2020), hierarchical Gaussian mixture models were used for combined inference in a simulated population of structures, for the purpose of damage identification.

### v2

@@ Abstract
Structural health monitoring (SHM) strategies seek to evaluate, predict, and maintain structural integrity, to improve the safety and design service life of structures in operation. Many of these strategies involve monitoring changes in structural dynamics, as damage can affect modal properties and present as changes in the characteristics of the resonance peaks of the frequency response function (FRF). While recent advances have improved the safety and reliability of structures, a number of challenges remain, impeding the practical implementation and generalisation of these systems. Like damage, benign variations, such as those caused by changes in temperature or other environmental fluctuations, can affect dynamic properties, making it difficult to distinguish between damage and normal operating conditions. In addition, newly-deployed structures can have insufficient data to describe the normal operating conditions (i.e., data scarcity), which can impair the development of data-based prediction models. Another common challenge is data loss (i.e., data sparsity), which may result from transmission issues, sensor failure, a sample-rate mismatch between sensors, and other causes. Missing data in the time domain will result in decreased resolution in the frequency domain, which can impair dynamic characterisation. For situations that may benefit from information sharing among datasets, e.g., population-based SHM of similar structures, the hierarchical Bayesian approach provides a useful modelling structure. Hierarchical Bayesian models learn statistical distributions at the population (or parent) and the domain levels simultaneously, to bolster statistical strength among the parameters. As a result, variance is reduced among the parameter estimates, particularly when data are limited. In this paper, a combined probabilistic FRF model is developed for a small population of nominally-identical helicopter blades, using a hierarchical Bayesian structure, to support information transfer in the context of sparse data. The modelling approach is also demonstrated in a traditional SHM context, for a single helicopter blade exposed to varying temperatures, to show how the inclusion of physics-based knowledge can improve generalisation beyond the training data, in the context of scarce data. These models address critical challenges in SHM, by accommodating benign variations that present as differences in the underlying dynamics, while also considering (and utilising), the similarities among the domains.
@@
Structural health monitoring (SHM); Population-based SHM (PBSHM); Hierarchical Bayes; Multilevel models; Uncertainty quantification; Repeatability
journal: Mechanical Systems and Signal Processing
@@ 1 Introduction
The current work is focussed on developing vibration-based structural health monitoring (SHM) strategies, that utilise data sharing, domain expertise, and/or physics-based knowledge, to enable information transfer among similar domains. Specifically, this work has involved the development of probabilistic frequency response function (FRF) models, using data collected from healthy, full-scale composite helicopter blades, with consideration for the similarities and differences among the data, and the resulting effects on the characteristics of the resonance peaks. Benign variations confounding monitoring systems is one of the major challenges in SHM, as even among nominally-identical structures, variations caused by manufacturing differences, ageing parts, and changes in testing conditions can introduce uncertainty in the underlying dynamics. One such example of the effects of manufacturing differences on dynamic properties was presented in P. Cawley et al. (1985), where Cawley et al. performed vibration testing on filament-wound carbon fibre-reinforced plastic (CFRP) tubes. Of the 18 tested, six tubes were considered ‘normal’ (i.e., having the same microstructure, with a ±$\pm 45^{\circ}$ fibre winding angle) P. Cawley et al. (1985). They found that the first and second natural frequencies of the ‘normal’ tubes varied by as much as 4% P. Cawley et al. (1985). When the remaining 12 tubes were also considered (which had intentional defects in their microstructure, such as slightly misaligned fibres, changes in volume fraction, etc.), the natural frequencies varied by as much as 18% P. Cawley et al. (1985). Likewise, changes in bolt tightness can similarly cause variations in natural frequency. Zhu et al. W.D. Zhu and He (2013) performed modal testing on an aluminium three-bay space-frame structure with bolted joints, before and after manually loosening several bolts to hand tight. They found that natural frequencies changed by as much as 8%, depending on the mode considered W.D. Zhu and He (2013). Global variations, such as changes in ambient temperature, can also affect modal properties. For example, increased temperature may reduce stiffness (depending on the specific material properties of the structure, and the duration/temperature of exposure), which may decrease natural frequency. Colakoglu M. Colakoglu (2008) tested a polyethylene fibre composite beam from ${}^{\circ}$ to ${}^{\circ}$, and found that the first natural frequency decreased by 12% over the measured temperature range. In addition, for the same polyethylene beam, Colakoglu M. Colakoglu (2008) found that modal damping increased with rising temperature. They concluded that the relationships between temperature and natural frequency, and between temperature and damping, appeared to be functional and monotonic for the polyethylene beam, over the temperature range considered M. Colakoglu (2008). Similar effects of temperature on stiffness/natural frequency and damping have been noted for a magnesium alloy H. Watanabe et al. (2004) and a composite honeycomb structure Y. Bai et al. (2018). Accounting for these benign fluctuations is important for the practical implementation and generalisation of SHM technologies, as features commonly used for damage identification may be sensitive to harmless changes as well as damage Worden et al. (2002); Sohn (2007); Alampalli (1998); Cawley (1997). (Indeed, it is well known that structural damage can reduce stiffness, often manifesting as a reduction in natural frequency. Benign variations can then either mimic or mask damage, depending on whether they exhibit a stiffening or softening effect Alampalli (1998); Cawley (1997); Sohn (2007). Similar effects have been noted with damping, and multiple studies have shown that damping tends to increase with damage, particularly with crack growth Cao et al. (2017).) Data scarcity and sparsity present additional challenges for SHM systems that rely on machine learning. Data scarcity refers to incomplete information regarding the damage- or normal-condition states of structures, particularly those newly in operation, and can impair model training and development. Likewise, sensing networks are prone to data loss (causing sparse data), because of sensor failure caused by harsh environmental conditions or insufficient maintenance. Transmission issues make wireless-sensing networks particularly susceptible to loss, and can be caused by large transmission distances between the sensors and base station J. Pei et al. (2005), software/hardware problems J. Meyer et al. (2010), and other issues such as weather changes, interference from nearby devices, or installation difficulties Y. Bao et al. (2013). Furthermore, modern systems that produce large amounts of high-resolution data can suffer losses resulting from a data transfer bottleneck A. Malekloo et al. (2022). Significant losses higher than 30% have been reported J. Meyer et al. (2010); Y. Bao et al. (2013); J. Lynch and Loh (2006), and a 0.38% data loss was found to have similar effects on power spectral density (PSD) as 5% additive noise T. Nagayama et al. (2007). Differences in sample rate among sensors could have a similar presentation, with the data captured at a lower rate seemingly missing data relative to that captured at a higher rate. Population-based SHM (PBSHM) addresses these challenges, by supporting knowledge transfer (e.g., damage or normal-condition states) within a population of structures, so that data-rich members can support those with less information. PBSHM strategies differ according to the similarity of the population members. Heterogeneous populations Gosliga et al. (2021); P.A. Gardner et al. (2021) refer to those comprised of highly disparate members, such as different suspension bridge designs, and may require extensive processing such as domain adaptation prior to information sharing. The current work considers a homogeneous population, comprised of nominally-identical members L.A. Bull et al. (2021), where variabilities in their dynamics result from minor discrepancies in geometry, boundary conditions, and environmental/operational changes. Homogeneous populations can be represented using a general model, called a population form, which attempts to capture the ‘essential nature’ of the population and benign variations among the members L.A. Bull et al. (2021). The form was first introduced in L.A. Bull et al. (2021), where a conventional or single Gaussian process (GP) was applied to frequency response functions (FRFs), to develop a representation for a nominally-identical population of eight degree-of-freedom (DOF) systems. Then, to accommodate greater differences among the nominally-identical members, an overlapping mixture of Gaussian processes (OMGP) Lázaro-Gredilla et al. (2012), was used in L.A. Bull et al. (2021, 2021), to infer multivalued wind-turbine power-curve data, with unsupervised categorisation of the data. The OMGP approach Lázaro-Gredilla et al. (2012) was again used in T.A. Dardeno et al. (2022) to develop a population form for real and imaginary FRFs, obtained from four nominally-identical, full-scale, composite helicopter blades. Recently, hierarchical modelling was used to improve the predictive capability of simulated L.A. Bull et al. (2022) and in-service L.A. Bull et al. (2023a) truck-fleet hazard models, and wind-turbine power curves L.A. Bull et al. (2023a). Specifically, the work presented in L.A. Bull et al. (2022, 2023a) used partial pooling, where data groups are treated as belonging to different domains, and each domain can be considered a realisation from global or population-level distributions. (This approach is in contrast to complete pooling, where all data are considered as belonging to a single domain, or no pooling, where each domain is treated as fully independent of the other domains). It was shown that when populations of structures were allowed to share correlated information, model uncertainty was reduced L.A. Bull et al. (2022, 2023a). In addition, domains with incomplete data were able to borrow statistical strength from data-rich groups L.A. Bull et al. (2022, 2023a). In L.A. Bull et al. (2023b), multilevel modelling with partial pooling was used to learn a 2D map of arrival times for waves propagating through a complex plate geometry, via GP regression, using a series of acoustic-emission experiments performed on the same plate but with differing experimental designs. Domain expertise regarding the positive gradient for the expected intertask functions was encoded into the model, via a linear mean function and appropriate priors L.A. Bull et al. (2023b); M.R. Jones et al. (2023). GP kernel hyperparameters were learnt at the domain level, to allow variation in the response surface among the different tests, and were correlated via a higher-level sampling distribution L.A. Bull et al. (2023b).
@@ 1.1 Research aims of the current work
The current work addresses critical SHM challenges, including uncertainties in the underlying dynamics of structures, and data-scarcity/sparsity issues, which can impair generalisation of SHM technologies. In line with the population forms developed in L.A. Bull et al. (2021, 2021); T.A. Dardeno et al. (2022); L.A. Bull et al. (2022, 2023a), this paper presents generalised, probabilistic FRF models that were developed for the helicopter blades in T.A. Dardeno et al. (2022), using a hierarchical Bayesian structure. Two case studies are presented. The first case used FRFs collected from all four blades, at ambient laboratory temperature, with variations among the blades resulting from manufacturing differences (e.g., small discrepancies in material properties and geometry) and boundary conditions. Limited training data that did not fully characterise the resonance peaks were taken from two of the FRFs, while sufficient training data were taken from the remaining two FRFs, so that information could be shared with the data-poor domains via shared distributions over the parameters. This situation is representative of incomplete data in the time domain, which would reduce the number of spectral lines in the frequency domain. Independent models were generated for comparison, to visualise the variance reduction from the combined model. The second case addressed another challenge, typical in a traditional SHM context, where insufficient information regarding the normal operating condition of a structure can impede model generalisation. In this case, vibration data were collected from a single helicopter blade, at various temperatures, in an environmental chamber. A probabilistic FRF model was again developed using a hierarchical approach with partial pooling, except, for this model, each domain represented a different temperature state. Functional relationships between temperature and natural frequency, and between temperature and damping, were approximated via Taylor series expansion, and polynomial coefficients were learnt at the ‘population’ level. A subset of temperature-varied FRFs were used to train the model. Higher-level distributions over the modal parameters, and the polynomial coefficients, were then used to generalise to temperature-varied FRFs not used to train the model. Model accuracy was evaluated by comparing the results to FRFs computed via measured vibration data.
@@ 2 Related work
Most literature related to population-level monitoring of engineering structures (or, in a more general SHM context, group-level monitoring of datasets), is focussed on transfer learning, which aims to improve predictions in a target domain given a source domain with more complete information. Some cases have involved fine-tuning the classifiers and weights of a pre-trained convolutional neural network (CNN) according to a new dataset. For example, CNNs have been used to detect cracks, as in R-T. Wu et al. (2019); Dorafshan et al. (2018); Gao and K.M. Mosalam (2018); Jang et al. (2019), and other defects (e.g., corrosion, staining), as in R-T. Wu et al. (2019); Perez et al. (2019). Other cases have used domain adaptation (DA), a subcategory of transfer learning, where the source and target domains have the same feature space, and the target domain is mapped onto a shared space. Predictions are then made from a single model. For example, domain-adversarial neural networks have been implemented for condition monitoring of a fleet of power plants Michau and Fink (2019), and for fault detection in a gearbox and three-story structure Ozdagli and Koutsoukos (2020). Population-level modelling can also be viewed in the context of multitask learning (MTL), where multiple tasks are solved simultaneously, while considering similarities (and differences) across tasks. In MTL, a combined inference allows sharing of correlated information among domain-specific models, to improve accuracy in data-poor domains L.A. Bull et al. (2023a); Sun et al. (2021). In the context of modelling engineering infrastructure, MTL has been used as part of a multioutput Gaussian process (GP) regression to learn correlations between data obtained at adjacent sensors from the Canton Tower, to reconstruct missing information H-P. Wan and Y-Q. Ni (2019). (Note that using MTL in this context differs from the current work. In H-P. Wan and Y-Q. Ni (2019), missing time-domain data from temperature and acceleration sensors were reconstructed using MTL, and all data were sourced from a single structure. The current work considers frequency-domain data from multiple structures, as in the first case, and incorporates the relationships between temperature and domain-specific tasks into the model structure, as in the second case.) A similar approach was used in Y. Li et al. (2021), where GP regression assisted with missing data recovery from a faulty sensor, by capturing the correlation among the remaining sensors on a hydroelectric dam. Likewise, in Seshadri et al. (2022), GPs were used to transfer information between spatial temperature or pressure profiles at axial stations within an aeroengine. Hierarchical Bayesian modelling is an MTL framework, where (lower-level) domain-specific tasks are correlated by conditioning on (higher or population-level), shared variables. An early introduction of these models to SHM was presented in G.C. Ballesteros et al. (2014), where a hierarchical Bayesian framework was proposed for accommodating variability in the elastic modulus of simulated bar structures, and in I. Behmanesh et al. (2015), which involved the model updating and uncertainty quantification of a numerical three-story shear-building model in the presence of environmental variations. Likewise, in Huang and J.L. Beck (2015); Huang et al. (2019), a multilevel hierarchical Bayesian model was developed to identify damage on a single structure by inferring stiffness loss from changes in modal parameters, given noisy, incomplete modal data. Modal parameters were learnt for different damage scenarios using a series of coupled linear regressions to approximate the eigenvalue problem Huang and J.L. Beck (2015); Huang et al. (2019). In Huang et al. (2019), the similarities between acceleration responses at adjacent sensors were exploited for data-loss recovery, by modelling the measured signal as a linear combination of basis functions. Both cases presented in Huang and J.L. Beck (2015); Huang et al. (2019) considered data from a single structure. (Note that the approaches presented in G.C. Ballesteros et al. (2014); I. Behmanesh et al. (2015); Huang and J.L. Beck (2015); Huang et al. (2019) differ from the current work. The work presented in the current paper, in the first case study, models the entire (band-limited) FRF from multiple structures, using a likelihood function based on modal parameters. The second case models the entire FRF using measurements at various temperatures from a single structure, and incorporates functional relationships between temperature and the modal parameters, enabling predictions at ‘unseen’ temperature conditions.) Recently, A. Poblete et al. (2022) used hierarchical Bayesian modelling to identify the electromechanical properties of energy harvesters, via individual and combined inference of experimentally-obtained FRFs. Population-level parameters were compared for partial- and complete-pooling approaches, successfully demonstrating that partial pooling results in improved estimates of population-level distributions compared to a classical Bayesian approach. The current work also utilises a hierarchical approach to probabilistic FRF modelling, but instead, the focus is on knowledge transfer - whereby one or more members of a population or group have markedly fewer data. The current work also increases the complexity of the population (parameter) models, to move beyond exchangeable tasks, and predict the model parameters for previously-unobserved experiments. (In other words, the current work allows for the prediction of unseen models, as well as data.) Hierarchical Bayesian models were again used in D. Di Francesco et al. (2021), to estimate corrosion rates via data from multiple sensors on the same test structure, to support decision-making in the absence of complete data. In Papadimas and Dodwell (2021) hierarchical Bayesian models were developed using strain measurements from multiple tensile tests to inform the material properties of the samples, in a manner that considered the inherent variability in material properties among the samples and uncertainties related to experimental repeatability. Likewise, in Dhada et al. (2020), hierarchical Gaussian mixture models were used for combined inference in a simulated population of structures, for the purpose of damage identification. Another recent development was presented in O. Sedehi et al. (2023), where a physics-informed GP model was developed using measured time-domain vibration data, with accommodation for temporal variability accomplished by partitioning the data into multiple subsegments, and applying a hierarchical Bayesian approach.